OpenStudy (anonymous):

Evaluate L^-1 {(s^2 - 29s + 5) / (s-4)^ (s^2 +3)} inverse laplace transform

OpenStudy (anonymous):

Is this what you want to find the inverse Lalplace Transform of \[ \frac{s^2-29 s+5}{(s-4)^2 \left(s^2+3\right)} \] I will suppose that that is it, since your are not online

OpenStudy (anonymous):

First use partial fraction \[ \frac{s^2-29 s+5}{(s-4)^2 \left(s^2+3\right)}=\frac{2-s}{s^2+3}+\frac{1}{s-4}-\frac{5}{(s-4)^2}=\\ \frac{2}{s^2+3}- \frac{s}{s^2+3}+\frac{1}{s-4}-\frac{5}{(s-4)^2} \]

OpenStudy (anonymous):

\[ \mathcal{L}_s^{-1}\left[\frac{2}{s^2+3}\right](t)=\frac{2 \sin \left(\sqrt{3} t\right)}{\sqrt{3}} \]

OpenStudy (anonymous):

\[ \mathcal{L}_s^{-1}\left[-\frac{s}{s^2+3}\right](t)=-\cos \left(\sqrt{3} t\right) \]

OpenStudy (anonymous):

\[ \mathcal{L}_s^{-1}\left[\frac{1}{s-4}\right](t)=e^{4 t} \]

OpenStudy (anonymous):

\[ \mathcal{L}_s^{-1}\left[-\frac{5}{(s-4)^2}\right](t)=-5 e^{4 t} t \]

OpenStudy (alekos):

good work elias, but where's nackulen?

OpenStudy (anonymous):

yeap that's the one..(referring to your first reply.)

OpenStudy (anonymous):

Did you understand the details?

OpenStudy (anonymous):

yes i do now. thank you so much.

OpenStudy (anonymous):

YW